OpenStudy (anonymous):

Check my answer? Indicated operations help please? Given the expression (6 + (sqrt)-100) - (3 + (sqrt)-49), perform the indicated operation and write the answer in the form of a + bi (A) 3 + 17i (B) -3 + 17i (C) 3 + 3i (D) -3 + 3i I'm leaning towards A or C but I'm not sure

OpenStudy (watchmath):

\((6+10i)-(3+7i)=3+3i\)

OpenStudy (danjs):

Recall a negative square root is not a real number: \[\sqrt{-1} = i\]

OpenStudy (danjs):

\[\sqrt{-100} = \sqrt{-1 * 100} = \sqrt{-1}\sqrt{100} = 10*i\]

OpenStudy (danjs):

similarly for the other root \[\sqrt{-49} = \sqrt{-1 * 49} = \sqrt{-1}\sqrt{49} = 7*i\]

OpenStudy (anonymous):

Hi Dan

OpenStudy (danjs):

hello, hows it goin

OpenStudy (anonymous):

I've come to find complex numbers more and more silly. Anyway, @JoeJoldin needs to respond

OpenStudy (danjs):

(6 + 10i) - (3 + 7i) 6 - 3 + 10i - 7i 3 +(10-7)i 3 + 3i

OpenStudy (danjs):

you understand all that ?

OpenStudy (danjs):

@JoeJoldin

OpenStudy (anonymous):

yes thank you

OpenStudy (danjs):

cool your welcome